# Strong solutions for a class of stochastic thermo-magneto-hydrodynamic-type systems with multiplicative noise

### v1

@@ Abstract.
We study the existence and uniqueness of strong solutions, in the sense of PDEs and probability, for a broad class of nonlinear stochastic partial differential equations (SPDEs) on a bounded domain $\mathscr{O}\subset\mathbb{R}^{d}$ ($d\leq 3$), perturbed by spatially correlated multiplicative noise. Our framework applies to many physically relevant systems, including stochastic convective Brinkman–Forchheimer equations, stochastic magnetohydrodynamics (MHD), stochastic Bénard convection in porous media, stochastic convective dynamo models, and stochastic magneto-micropolar fluids, among others. The analysis relies on Galerkin approximations and compactness arguments. Up to a suitable stopping time, we derive strong moment bounds and verify a Cauchy property for the approximate solutions, in the absence of any inherent cancellation structure. By applying a Gronwall-type lemma for stochastic processes, we establish the existence and uniqueness of maximal strong pathwise solutions, which are global in two spatial dimensions. These results provide a unified treatment of a wide class of nonlinear stochastic thermo-magneto-fluid models in the literature.
@@ 1. Introduction
This paper investigates the existence and uniqueness of strong solutions (in the sense of PDEs and probability) for a class of abstract nonlinear stochastic partial differential equations (SPDEs) defined on a bounded domain $\mathscr{O}\subset\mathbb{R}^{d}$, where $d=2,3$. These problems are motivated by thermally-coupled models in magneto-hydrodynamics, which describe the interplay between thermal, magnetic, and fluid dynamic phenomena under stochastic influences. We consider stochastic evolution equations of the form
$\partial_{t}\Phi+\mathcal{A}\Phi+\mathcal{B}(\Phi,\Phi)+\mathcal{R}(\Phi)$ $=g(\Phi)\dot{W},$ (1.1)
@@
which describes the time evolution of a state vector $\Phi$. Here, $\mathcal{A}$ is a linear operator representing dissipative effects, $\mathcal{B}$ is a bilinear form capturing convective nonlinearities, $\mathcal{R}$ is a nonlinear operator accounting for other reactive physical effects. The term $g(\Phi)\dot{W}$ represents a multiplicative noise, which is white in time and spatially correlated, modelling random forcing that depends on the state of the system. We impose a set of conditions on the maps $\mathcal{A}$, $\mathcal{B}$, $\mathcal{R}$, and $g$ to cover a range of physically relevant thermo-magneto-hydrodynamical systems.
While the existence of martingale weak or pathwise weak solutions to SPDEs arising from specific magneto-hydrodynamical models, including the stochastic Navier–Stokes [6, 19], Boussinesq [15, 52], or MHD equations [38, 43], has been addressed in a number of separate works, these studies typically analyse each equation in isolation. In contrast, this paper develops an abstract setting that encompasses a wide range of thermally-coupled stochastic magneto-hydrodynamical-type systems. Beyond unifying the treatment of previously studied models, our formulation also captures several important stochastic nonlinear systems that have not, to our knowledge, been rigorously analysed before in either the weak or strong sense.
The existence of pathwise weak solutions for a class of SPDEs arising in fluid dynamics is established in [12]. While our work is related to that result, the contributions here are fundamentally different in both analytical strength and scope. Specifically, [12] proves the existence of pathwise solutions that are weak in the analytical sense, whereas we establish the existence of pathwise solutions that are analytically strong. In addition, our framework accommodates equations with more general nonlinearities, extending beyond the convective-type nonlinearities treated in other works. Examples of models covered by our setting, but not necessarily by that in [12], are given in Section 3, where some of the corresponding well-posedness results appear to be new.
@@
Our approach also departs significantly from that in [12]. Although both proofs employ Galerkin approximations and compactness arguments, our analysis requires substantially more delicate estimates. In particular, up to a suitable stopping time, we must derive strong moment bounds and verify a Cauchy property for the approximate solutions in the absence of any inherent cancellation structure. Finally, by applying a Gronwall-type lemma for stochastic processes from [24], we conclude the existence and uniqueness of maximal strong pathwise solutions, which are global in the case $d=2$.
@@
  • Section 5 extends the local strong solution to a maximal (random) time interval and shows that this maximal strong solution is global when $d=2$.
@@
Our main results are the existence of a unique maximal strong pathwise solution for $d\leq 3$ (Theorem 5.2) and the existence of a global strong pathwise solution for $d=2$ (Theorem 5.3).
@@ 2.1. Functional setting
Let $\mathscr{O}\subset\mathbb{R}^{d}$, for $d=2,3$, be an open bounded domain with $C^{2}$-smooth boundary or a convex Lipschitz domain. Let $(\mathcal{H},\left|{\cdot}\right|)$ be a separable Hilbert space. Let $\mathcal{A}$ be an unbounded, self-adjoint, positive linear operator on $\mathcal{H}$, defined on a domain $\mathrm{D}(\mathcal{A})\subset\mathcal{H}$, such that $\mathcal{A}$ is a bijection onto $\mathcal{H}$ with a compact inverse. Consequently, there exists an orthonormal basis $\{e_{k}\}_{k\in\mathbb{N}}$ for $\mathcal{H}$ consisting of eigenfunctions of $\mathcal{A}$. The associated positive eigenvalues $\{\lambda_{k}\}_{k\in\mathbb{N}}$ form an unbounded increasing sequence. Given $\alpha>0$, let
@@ 2.2. Assumptions
  • $\mathcal{R}:\mathcal{V}\to\mathcal{V}^{\prime}$ is a continuous map defined by $\mathcal{R}(v)=\mathcal{S}(v)+\mathcal{F}(v)$ satisfying $\mathcal{F}(v)\in\mathcal{H}$, such that for any $v\in\mathcal{V}$,
@@
  • There exists $C>0$ such that for any $v_{1}\in\mathrm{D}(\mathcal{A})$ and $v_{2}\in\mathcal{V}$,
    $\left|{\left\langle\mathcal{R}(v_{1}),v_{2}\right\rangle}\right|\leq\begin{cases}C\left(1+\left|{v_{1}}\right|^{\frac{3}{2}}\left|{\mathcal{A}v_{1}}\right|^{\frac{1}{2}}\right)\left\|{v_{1}}\right\|\left|{v_{2}}\right|,&\text{if $d=2$},\\[4.30554pt] C\left(1+\left|{v_{1}}\right|^{\frac{1}{2}}\left\|{v_{1}}\right\|\left|{\mathcal{A}v_{1}}\right|^{\frac{1}{2}}\right)\left\|{v_{1}}\right\|\left|{v_{2}}\right|,&\text{if $d\leq 3$}.\end{cases}$ (2.7)
  • There exists $C>0$ such that for any $v_{1},v_{2}\in\mathcal{V}$ and $v_{3}\in\mathcal{V}$,
    $\left|{\left\langle\mathcal{R}(v_{1})-\mathcal{R}(v_{2}),v_{3}\right\rangle}\right|\leq C\left(1+\left\|{v_{1}}\right\|^{2}+\left\|{v_{2}}\right\|^{2}\right)\left\|{v_{1}-v_{2}}\right\|\left|{v_{3}}\right|.$ (2.8)
@@
    $\left|{\mathcal{R}(v)}\right|^{2}$ $\leq C\left(1+\left\|{v}\right\|^{4}\right)\left\|{v}\right\|^{2},$ (2.9)
@@ 3. Physical models
In this section, we present several physically relevant models that fall within the scope of our framework. We note, however, that our assumptions are not limited to these thermo-magneto-fluid models and can accommodate a broader class of nonlinear stochastic PDEs arising in various contexts.
@@ 3.1. Stochastic convective Brinkman–Forchheimer equations
The existence and uniqueness of probabilistically and analytically strong solutions to (3.1) in the torus (or a domain with periodic boundary conditions) are established in [30, 37]. For the case of Dirichlet boundary condition, a remark is made in [34, 37] that the arguments presented in [23, 35] contain a gap stemming from the use of an incorrect identity in various parts of the proof, which invalidates the corresponding existence result. In this work, we correct these mistakes and establish the existence of pathwise strong solutions for $r\in[2,3]$, including the case of Dirichlet boundary condition. When $d=2$, the solutions are shown to exist globally in time. Our results also extend to the case of Navier-type (or free) boundary conditions [4, 56].
@@
By means of the Helmholtz–Leray projector $\mathbb{P}:L^{2}(\mathscr{O})\to\mathcal{H}$, noting that $\mathbb{P}v=v$ and $\mathbb{P}[g(v)\dot{W}]=\sum_{k=1}^{\infty}\mathbb{P}[g_{k}(v)]\dot{\beta}_{k}$ for $v\in\mathcal{H}$ due to (2.11) and (2.12), we can project (3.1) onto the space of divergence-free vector fields to formulate problem (3.1) in the form (1.1), with
@@
The map $\mathcal{R}$ clearly satisfies (2.6) with $\mathcal{F}(v)=0$. It remains to show that the map satisfies (2.7), (2.8), (2.9), and (2.10). For ease of presentation, we assume $r=3$, noting that similar argument works for $r\leq 3$. First, we show (2.7). By the Hölder inequality (with exponents $p=q=4$ when $d=2$, and $p=6,q=3$ when $d=3$) and the Gagliardo–Nirenberg inequalities, we have for any $v_{1},v_{2}\in\mathcal{V}$,
$\left|{\left\langle\mathcal{R}(v_{1}),v_{2}\right\rangle}\right|$ $\leq C\left(\left|{v_{1}}\right|+\left|{v_{1}}\right|_{L^{\infty}}\left|{v_{1}}\right|_{L^{p}}\left|{v_{1}}\right|_{L^{q}}\right)\left|{v_{2}}\right|_{L^{2}}$
@@
which implies (2.7). By a similar argument, for $d\leq 3$ and any $v_{1},v_{2}\in\mathcal{V}$ and $v_{3}\in\mathcal{H}$, we obtain
$\left|{\left\langle\mathcal{R}(v_{1})-\mathcal{R}(v_{2}),v_{3}\right\rangle}\right|$ $\leq\left(\left|{v_{1}}\right|_{L^{6}}^{2}+\left|{v_{2}}\right|_{L^{6}}\left|{v_{1}+v_{2}}\right|_{L^{6}}\right)\left|{v_{1}-v_{2}}\right|_{L^{6}}\left|{v_{3}}\right|_{L^{2}}$
$\leq C\left(1+\left\|{v_{1}}\right\|^{2}+\left\|{v_{2}}\right\|^{2}\right)\left\|{v_{1}-v_{2}}\right\|\left|{v_{3}}\right|$
by the Sobolev embedding $\mathcal{V}\hookrightarrow L^{6}$ and Young’s inequality. This shows (2.8). Similarly, for any $v\in\mathcal{V}$, we have
thus showing (2.9). Finally, we will show (2.10). To this end, note that $\left|{v}\right|^{2}v\in H^{1}(\mathscr{O})$ for any $v\in\mathrm{D}(\mathcal{A})$ since by Hölder’s inequality and the Sobolev embedding,

### v2

@@ Abstract.
We establish the existence and uniqueness of strong solutions, in both the PDE and probabilistic sense, for a broad class of nonlinear stochastic partial differential equations (SPDEs) on a bounded domain $\mathscr{O}\subset\mathbb{R}^{d}$ ($d=2,3$), driven by multiplicative Gaussian noise. The solutions are global in time for $d=2$. This theory simultaneously covers several physically relevant systems, including stochastic convective Brinkman–Forchheimer equations, stochastic magnetohydrodynamics (MHD), stochastic Bénard convection in porous media, stochastic convective dynamo system, stochastic thermo-magneto-micropolar fluids, and stochastic diffusive tropical climate model, for which previous results only provide analytically weak martingale or pathwise solutions. The proof relies on Galerkin approximation and compactness argument. Up to a suitable stopping time, we derive strong moment bounds and verify a Cauchy property for the approximate solutions, in the absence of any inherent cancellation structure. By applying a Gronwall-type lemma for stochastic processes, we establish the existence and uniqueness of maximal strong pathwise solutions, which are global in two spatial dimensions.
@@ 1. Introduction
This paper investigates the existence and uniqueness of strong solutions, in the sense of PDEs and probability, for a class of abstract nonlinear stochastic partial differential equations (SPDEs) defined on a bounded domain $\mathscr{O}\subset\mathbb{R}^{d}$, where $d=2,3$. These problems are motivated by thermally-coupled models in magneto-hydrodynamics (MHD), which describe the interplay between thermal, magnetic, and fluid dynamic phenomena under stochastic influences. The problem takes this general form:
${\frac{\mathrm{d}\Phi}{\mathrm{d}t}}+\mathcal{A}\Phi+\mathcal{B}(\Phi,\Phi)+\mathcal{R}(\Phi)$ $=g(\Phi)\dot{W},$ (1.1)
@@
which describes the time evolution of a state vector $\Phi$. Here, $\mathcal{A}$ is a linear operator representing dissipative effects, $\mathcal{B}$ is a bilinear form capturing convective nonlinearities, and $\mathcal{R}$ is a nonlinear operator accounting for other reactive physical effects. The term $g(\Phi)\dot{W}$ represents a multiplicative Gaussian noise modelling random forcing that depends on the state of the system. To cover a range of physically relevant thermo-magneto-hydrodynamical systems, a set of assumptions on the maps $\mathcal{A}$, $\mathcal{B}$, $\mathcal{R}$, and $g$ will be imposed later. These assumptions ensure that (1.1) encompasses several physically relevant models, such as:
  1. (1)
    the stochastic convective Brinkman–Forchheimer equations [8, 22, 25, 30, 37];
  2. (2)
    the stochastic magnetohydrodynamic (MHD) equations [38, 42, 43, 45, 52];
  3. (3)
    the stochastic Boussinesq model for Bénard convection [11, 14, 46, 52];
  4. (4)
    the stochastic convective dynamo system [9, 10, 47];
  5. (5)
    the stochastic thermo–magneto–micropolar fluid systems [28, 29, 48, 50, 51]; and
  6. (6)
    the stochastic diffusive tropical climate model [19, 33, 53, 54],
among others. These problems will be elaborated further in Section 3.
The works cited above address a broad range of topics, such as the physical derivations of the models, various notions of well-posedness, existence of invariant measures, asymptotic behaviours, and large deviation principles. In what follows, we restrict our attention to the analytical studies concerning the existence of solutions. In particular, for many of the problems listed above, the existence of analytically weak martingale or analytically weak pathwise solutions has been studied in separate works, which we now briefly summarise:
  1. (1)
    Stochastic convective Brinkman–Forchheimer equations: Global martingale weak and pathwise weak solutions are shown to exist in [22] for $d=2,3$ under Dirichlet boundary condition.
  2. (2)
    Stochastic MHD equations: For $d=2,3$, the existence of global martingale weak solutions is established in [43, 52] for bounded domains, and in [38] for more general Poincaré domains. Pathwise uniqueness is also shown for $d=2$, which implies the existence of pathwise weak solutions.
  3. (3)
    Stochastic Boussinesq model for Bénard convection: Local pathwise weak solutions exist in bounded domains for $d=2,3$, as shown in [11], and these solutions are global for $d=2$.
  4. (4)
    Stochastic convective dynamo system: To the best of our knowledge, the existence of solutions for this system has not been established yet.
  5. (5)
    Stochastic thermo–magneto–micropolar fluid: Global martingale weak solutions exist for $d=2,3$, as proven in [50, 51].
  6. (6)
    Stochastic diffusive tropical climate model: Local pathwise weak solutions exist for two-dimensional unbounded domains, as shown in [54].
In contrast, the existence of analytically strong pathwise solutions for these problems in bounded domains with physically relevant boundary conditions is currently unknown. It is worth noting that for the stochastic convective Brinkman–Forchheimer equations, the existence and uniqueness of probabilistically and analytically strong solutions on a torus (or domains with periodic boundary conditions) are established in [30, 37]. For the case of Dirichlet boundary condition, [34, 37] remark that the arguments in [23, 35] contain a gap due to the use of some incorrect identities in several parts of the proof. Moreover, the authors implicitly assume throughout the paper that the Leray projector commutes with the Laplacian, which does not hold under Dirichlet boundary conditions. Consequently, the corresponding existence result for the Dirichlet case is invalid.
The primary focus of this work is the rigorous proof of the existence of analytically strong pathwise solutions for a wide class of thermally-coupled stochastic MHD-type systems for $d=2,3$, with solutions being global in time for $d=2$. To the best of our knowledge, this provides the first strong solution theory for several of these stochastic nonlinear systems, which were previously understood only in the weak sense or not at all. In addition, we develop an abstract framework that unifies the treatment of these models, encompassing both previously studied systems and new physically relevant stochastic nonlinear systems that have not been analysed rigorously before.
@@
The proofs are based on Galerkin approximations combined with compactness arguments, but the analysis requires substantially more delicate estimates than in previous works. In particular, following [24], we derive strong moment bounds and establish a Cauchy property for the approximate solutions up to a suitable stopping time, despite the absence of any inherent cancellation structure in the equations. The final step employs a Gronwall-type lemma for stochastic processes, which allows us to conclude the existence and uniqueness of maximal strong pathwise solutions, with global-in-time existence guaranteed for $d=2$.
@@
  • Section 5 extends the local strong solution to a maximal (random) time interval for $d=2,3$, and shows that this maximal strong solution is global when $d=2$.
@@
Our main results are the existence of a unique maximal strong pathwise solution for $d=2,3$ (Theorem 5.2) and the existence of a global strong pathwise solution for $d=2$ (Theorem 5.3).
@@ 2. Preliminaries
2.1. Notations
We begin by defining some notations used in this paper. For $p\in[1,\infty]$ and $k\in\mathbb{N}$, the function spaces $L^{p}$ and $W^{k,p}$ denote, respectively, the Lebesgue space of $p$-th integrable functions and the Sobolev space of $L^{p}$ functions whose weak derivatives up to order $k$ also belong to $L^{p}$, respectively. As usual, we set $H^{k}:=W^{k,2}$.
If $X$ is a Banach space, $L^{p}(0,T;X)$ and $W^{k,p}(0,T;X)$ denote, respectively, the usual Lebesgue and Sobolev spaces of strongly measurable functions on $(0,T)$ taking values in $X$. Similarly, the space $L^{p}(\Omega;X)$ denotes the space of $X$-valued strongly measurable random variables with finite $p$-th moment, where $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space. The space $C([0,T];X)$ and $C_{\mathrm{w}}([0,T];X)$ denote the spaces of functions on $[0,T]$ taking values in $X$ that are continuous with respect to the strong and weak topologies of $X$, respectively.
In the analysis, the constant $C$ denotes a generic constant which may take different values at different occurrences. If the dependence of $C$ on a variable, e.g. $s$, is highlighted, we will write $C_{s}$.
@@ 2.2. Functional setting
Let $(\mathcal{H},\left|{\cdot}\right|)$ be a separable Hilbert space and let $\mathcal{A}$ be an unbounded, self-adjoint, positive linear operator on $\mathcal{H}$, defined on a domain $\mathrm{D}(\mathcal{A})\subset\mathcal{H}$, such that $\mathcal{A}$ is a bijection onto $\mathcal{H}$ with a compact inverse. Consequently, there exists an orthonormal basis $\{e_{k}\}_{k\in\mathbb{N}}$ for $\mathcal{H}$ consisting of eigenfunctions of $\mathcal{A}$. The associated positive eigenvalues $\{\lambda_{k}\}_{k\in\mathbb{N}}$ form an unbounded increasing sequence. Given $\alpha>0$, let
@@ 2.3. Assumptions
  • $\mathcal{R}:\mathcal{V}\to\mathcal{H}$ is a continuous map defined by $\mathcal{R}=\mathcal{S}+\mathcal{F}$ where $\mathcal{S},\mathcal{F}:\mathcal{V}\to\mathcal{H}$ satisfy, for any $v\in\mathcal{V}$,
@@
  • There exists $C>0$ such that for any $v\in\mathcal{V}$,
    $\left|{\mathcal{R}(v)}\right|$ $\leq C\left(1+\left\|{v}\right\|^{2}\right)\left\|{v}\right\|.$ (2.7)
    Moreover, for any $v\in\mathrm{D}(\mathcal{A})$,
    $\left|{\mathcal{R}(v)}\right|\leq\begin{cases}C\left(1+\left|{v}\right|^{\frac{3}{2}}\left|{\mathcal{A}v}\right|^{\frac{1}{2}}\right)\left\|{v}\right\|,&\text{if $d=2$},\\[4.30554pt] C\left(1+\left|{v}\right|^{\frac{1}{2}}\left\|{v}\right\|\left|{\mathcal{A}v}\right|^{\frac{1}{2}}\right)\left\|{v}\right\|,&\text{if $d\leq 3$}.\end{cases}$ (2.8)
  • There exists $C>0$ such that for any $v_{1},v_{2}\in\mathcal{V}$
    $\left|{\mathcal{R}(v_{1})-\mathcal{R}(v_{2})}\right|_{\mathcal{H}}\leq C\left(1+\left\|{v_{1}}\right\|^{2}+\left\|{v_{2}}\right\|^{2}\right)\left\|{v_{1}-v_{2}}\right\|$ (2.9)
@@ 3. Physical models
We now present several physically relevant models that fall within the scope of our framework. We note, however, that our assumptions are not limited to these thermo-magneto-fluid models and can accommodate a broader class of nonlinear stochastic PDEs arising in various contexts.
In the following sections, $\mathscr{O}\subset\mathbb{R}^{d}$, for $d=2,3$, denotes an open bounded domain with $C^{2}$-smooth boundary or a convex Lipschitz domain.
@@ 3.1. Stochastic convective Brinkman–Forchheimer equations
The existence and uniqueness of probabilistically and analytically strong solutions to (3.1) on a torus (or a domain with periodic boundary conditions) are established in [30, 37]. For the case of Dirichlet boundary condition, a remark is made in [34, 37] that the arguments presented in [23, 35] contain a gap stemming from the use of an incorrect identity in various parts of the proof, which invalidates the corresponding existence result. In this work, we correct these mistakes and establish the existence of pathwise strong solutions for $r\in[2,3]$, including the case of Dirichlet boundary condition. When $d=2$, the solutions are shown to exist globally in time. Our results also extend to the case of Navier-type (or free) boundary conditions [4, 56].
@@
In the definition of $\mathcal{H}$ here and in what follows, the condition $\mathrm{div}\,u=0$ is understood in the sense of distributions, while the trace of the normal component of $u$ on the boundary is interpreted in a suitable generalised sense (see [44, Chapter I, Section 1.3] and [20, Section III.2]). By means of the Helmholtz–Leray projector $\mathbb{P}:L^{2}(\mathscr{O})\to\mathcal{H}$, noting that $\mathbb{P}v=v$ and $\mathbb{P}[g(v)\dot{W}]=\sum_{k=1}^{\infty}g_{k}(v)\dot{\beta}_{k}$ due to (2.11) and (2.12), we can project (3.1) onto the space of divergence-free vector fields to formulate problem (3.1) in the form (1.1), with
@@
The map $\mathcal{R}$ clearly satisfies (2.6) with $\mathcal{F}(v)=0$. It remains to show that the map satisfies (2.7), (2.8), (2.9), and (2.10). For ease of presentation, we assume $r=3$, noting that similar argument works for $r\leq 3$. First, for any $v\in\mathcal{V}$, we have by Hölder’s inequality and the Sobolev embedding,
thus showing (2.7). Next, we show (2.8). By the Hölder inequality (with exponents $p=q=4$ when $d=2$, and $p=6,q=3$ when $d=3$) and the Gagliardo–Nirenberg inequalities, we have for any $v_{1},v_{2}\in\mathcal{V}$,
$\left|{\left\langle\mathcal{R}(v_{1}),v_{2}\right\rangle}\right|$ $\leq C\left(\left|{v_{1}}\right|_{L^{2}}+\left|{v_{1}}\right|_{L^{\infty}}\left|{v_{1}}\right|_{L^{p}}\left|{v_{1}}\right|_{L^{q}}\right)\left|{v_{2}}\right|_{L^{2}}$
@@
as required. By a similar argument, for $d\leq 3$ and any $v_{1},v_{2}\in\mathcal{V}$ and $v_{3}\in\mathcal{H}$, we obtain
$\left|{\mathcal{R}(v_{1})-\mathcal{R}(v_{2})}\right|$ $\leq\left(\left|{v_{1}}\right|_{L^{6}}^{2}+\left|{v_{2}}\right|_{L^{6}}\left|{v_{1}+v_{2}}\right|_{L^{6}}\right)\left|{v_{1}-v_{2}}\right|_{L^{6}}$
$\leq C\left(1+\left\|{v_{1}}\right\|^{2}+\left\|{v_{2}}\right\|^{2}\right)\left\|{v_{1}-v_{2}}\right\|$
by the Sobolev embedding $\mathcal{V}\hookrightarrow L^{6}$ and Young’s inequality. This shows (2.9). Finally, we will show (2.10). To this end, note that $\left|{v}\right|^{2}v\in H^{1}(\mathscr{O})$ for any $v\in\mathrm{D}(\mathcal{A})$ since by Hölder’s inequality and the Sobolev embedding,